student174:

Type the correct answer in the box. Use numerals instead of words. If necessary, use / for the fraction bar. A rectangle is 12 feet long and 5 feet wide. If the length of the rectangle is increased by 25% and the width is decreased by 20%, what is the change in the area of the rectangle? The change in the area of the rectangle is square feet.

supie:

We have to calculate the area of ​​the two rectangles so, The formula: \[a = area\] \[l = length = 12ft\] \[w = width = 5ft\] \[a = l * w\] Now we change the values that are already recognized \[a = 12ft * 5ft\] \[a = 60ft²\] Now we do the same for the other triangle \[w = 5ft - (5ft * 20/100)\] \[w = 5ft - 1ft\] \[w = 4ft\] \[l = 12ft + (12ft * 25/100)\] \[l = 12ft + 3ft\] \[l = 15ft\] \[a2 = 15ft * 4ft\] \[a2 = 60ft²\] Now we subtract (a - a2) to calculate the change in the area \[a - a2 =\] \[60ft² - 60ft² = 0ft²\]